# LongEmbed: Extending Embedding Models for Long Context Retrieval

### v2

@@ Abstract
Embedding models play a pivot role in modern NLP applications such as IR and RAG. While the context limit of LLMs has been pushed beyond 1 million tokens, embedding models are still confined to a narrow context window not exceeding 8k tokens, refrained from application scenarios requiring long inputs such as legal contracts. This paper explores context window extension of existing embedding models, pushing the limit to 32k without requiring additional training. First, we examine the performance of current embedding models for long context retrieval on our newly constructed LongEmbed benchmark. LongEmbed comprises two synthetic tasks and four carefully chosen real-world tasks, featuring documents of varying length and dispersed target information. Benchmarking results underscore huge room for improvement in these models. Based on this, comprehensive experiments show that training-free context window extension strategies like position interpolation can effectively extend the context window of existing embedding models by several folds, regardless of their original context being 512 or beyond 4k. Furthermore, for models employing absolute position encoding (APE), we show the possibility of further fine-tuning to harvest notable performance gains while strictly preserving original behavior for short inputs. For models using rotary position embedding (RoPE), significant enhancements are observed when employing RoPE-specific methods, such as NTK and SelfExtend, indicating RoPE’s superiority over APE for context window extension. To facilitate future research, we release E5Base-4k and E5-RoPEBase, along with the LongEmbed benchmark.
Figure 1: (a) Overview of the LongEmbed benchmark. (b) Performance of current embedding models on passkey retrieval, with evaluation length ranging from 256 to 32,768.222For simplicity, we report results from the base versions of the included models by default. $\blacktriangle$ / $\blacklozenge$ denotes embedding models with 512 / $\geq$ 4k context. The greener   a cell is, the higher retrieval accuracy this model achieves on the corresponding evaluation length. (c) Effects of context window extension methods on E5, E5-RoPE, E5-Mistral, measured by improvements of Avg. Scores on LongEmbed. SE / NTK is short for SelfExtend / NTK-Aware Interpolation.
@@ 1 Introduction
Text embeddings are vector representations of natural language that encode its semantic information. They play a pivotal role in various natural language processing (NLP) tasks, including information retrieval (IR) and retrieval-augmented generation (RAG). However, embedding models for producing these vector representations still operates within a very narrow context window, typically 512 input tokens (Wang et al., 2022; Xiao et al., 2023; Ni et al., 2022), This narrow context window has greatly hindered their application in scenarios requiring long inputs, such as long wikipedias and meeting scripts (Saad-Falcon et al., 2024).
@@
For models utilizing RoPE (Su et al., 2021), substantial enhancements on LongEmbed are observed when employing methods that fully leverage RoPE’s advantages, such as NTK (Peng & Quesnelle, 2023) and SelfExtend (Jin et al., 2024). As illustrated in Figure 1(b) and 1(c), leveraging NTK extends the context window of E5-Mistral to 32k, achieving close-to-perfect accuracy on passkey retrieval and state-of-the-art performance on LongEmbed. Further, for fair comparison of APE / RoPE-based embedding models, we pre-train E5-RoPE following the training procedure and data of E5. Thorough comparison of E5 and E5-RoPE reveals the superiority of RoPE-based embedding models in context window extension.
To facilitate future research in long context embedding models, we release E5Base-4k, E5-RoPEBase, and the LongEmbed benchmarks. E5Base-4k is further fine-tuned on E5Base to support 4k context, while strictly preserving original behavior for inputs not exceeding 512 tokens. E5-RoPEBase follows the same training procedure as E5Base, except for the substitution of APE with RoPE. It is released to facilitate comparison between APE & RoPE-Based embedding models. Furthermore, we have integrated LongEmbed into MTEB (Muennighoff et al., 2023) to make evaluation more convenient.
@@ 2 Related Work
Text Embedding Models. Text embeddings are continuous, low-dimensional vector representations of text that encode semantic information, laying the foundation of numerous NLP applications. Early attempts on text embeddings includes latent semantic indexing (Deerwester et al., 1990) and weighted average of word embeddings (Mikolov et al., 2013). Modern embedding models (Wang et al., 2022; Xiao et al., 2023; Neelakantan et al., 2022) exploit supervision from labeled query-document pairs, adopting a multi-stage training paradigm, where they are first pre-trained on large-scale weakly-supervised text pairs using contrastive loss, then fine-tuned on small scale but high-quality datasets. More recently, Muennighoff et al. (2024) explores the combination of generative and embedding tasks on LLMs, introducing GritLM that harvests improvements in both aspects.
@@
Context Window Extension for Large Language Models. Due to the high cost of pre-training an LLM from scratch, there have been many efforts towards extending the context window of existing LLMs in a plug-and-play manner. We categorize these efforts as follows: 1) Divide-and-conquer, which involves segmenting long inputs into short chunks, processing each chunk with the model, and aggregating the results, as demonstrated by PCW (Ratner et al., 2023); 2) Position reorganization, which reorganizes position ids to boost length extrapolation, as exemplified by SelfExtend (Jin et al., 2024), DCA (An et al., 2024), and others; 3) Position interpolation, which introduces new position embeddings by interpolating existing ones, includes PI (Chen et al., 2023), NTK (Peng & Quesnelle, 2023), YaRN (Peng et al., 2023), and Resonance RoPE (Wang et al., 2024a). Our paper thoroughly investigates these three lines of methods on embedding models. We also acknowledge other efforts for extending the context window, such as prompt & KV compression (Jiang et al., 2023; Ge et al., 2023; Zhang et al., 2024a) and memory-based transformers (Wang et al., 2024b; Xiao et al., 2024). However, the former is not applicable for bidirectional attention, and the latter requires complex mechanisms for accessing encoded content, hence we do not experiment with these two categories.
In addition to their plug-and-play usability, further fine-tuning on top of these methods with long training samples has been proven to yield better performance  (Xiong et al., 2023; Fu et al., 2024; Zhang et al., 2024b; Yen et al., 2024). Addressing the overhead of training on long inputs and the scarcity of extremely long training data, a line of research investigates simulating long inputs within short context, including Randomized Positions (Ruoss et al., 2023), Positional Skip-wise (PoSE) training (Zhu et al., 2023). This paper also leverage these efforts to synthesize long training samples from the original training data, facilitating further fine-tuning on top of plug-and-play methods.
@@ 3 The LongEmbed benchmark
In this section, we first identify two limitations of existing retrieval benchmarks for evaluating long-context capabilities (Section 3.1). Then, we introduce the retrieval tasks adopted in our LongEmbed, including both synthetic ones (Section 3.2) and real ones (Section 3.3).
3.1 Examination of Existing Retrieval Benchmarks
There are mainly two desiderata for curating a benchmark for long context retrieval. First, the candidate documents should be long enough. Second, the target information to answer user query should be as uniformly distributed across the document as possible. This prevents embedding models from solely focusing on specific parts, such as the beginning (Coelho et al., 2024), to achieve unreasonably high scores. Based on these criteria, we evaluate existing benchmarks for text retrieval as follows:
Figure 2: Results of E5Base on 8 LoCo tasks.
BEIR Benchmark (Thakur et al., 2021) is a collection of 18 information retrieval datasets, ranging across ad-hoc web search, question answering, fact verification and duplicate question retrieval, etc. However, documents in this benchmark contains fewer than 300 words on average (See Table 5 in Appendix), making it unsuitable for measuring long context retrieval that usually involves documents of thousands or tens of thousands of words.
LoCo Benchmark (Saad-Falcon et al., 2024) consists 12 retrieval tasks that requires long context reasoning, spanning diverse domains such as law, science, finance, etc. However, we show that it still suffers from biased distribution of key information. Figure 2 presents results of E5Base on 8 LoCo tasks that are publicly available. With only 512 context length, E5Base achieves >85% nDCG scores on 3 out of 8 retrieval tasks. This severely biased distribution of target information undermines its ability to reflect model performance as context length increases.
3.2 Synthetic Tasks in LongEmbed
First, we tailor the passkey retrieval and needle-in-a-haystack retrieval task designed for LLMs to measure context length of embedding models as follows:
Personalized Passkey Retrieval. Passkey retrieval (Mohtashami & Jaggi, 2023) requires LLMs to recover a random passkey hidden within a long document comprising garbage information. For embedding models, we adopt the personalized passkey retrieval task proposed by Wang et al. (2023b), where each document contains a unique person name and his/her passkey at random position. The goal is to retrieve the document containing the given person’s passkey from all candidates documents.
Figure 3: Example for the passkey and needle test.
Needle-in-a-haystack Retrieval. While passkey retrieval surrounds key information with garbage sentences, needle-in-a-haystack retrieval (Kamradt, 2023) randomly inserts key information into an arbitrary position of a long essay, making the task more challenging. To tailor this task for embedding models, we instruct GPT-4 to generate 100 facts covering a variety of domains including physics, history, geometry, art, etc, and 100 queries correspondingly. The facts are treated as needles and randomly inserted into the PaulGrahamEssay to form 100 candidate documents. Our task is to correctly retrieve the document that contains corresponding needle given the query.
The advantage of synthetic data is that we can flexibly control context length and distribution of target information. For both tasks, we evaluate a broad context range of $\{0.25,0.5,1,2,4,8,16,32\}\times 1024$ tokens  [^3] Since token numbers vary w.r.t. tokenizers, we use a rough estimation that 1 token = 0.75 word, and constraint the word numbers to not exceed $\{0.25,0.5,1,2,4,8,16,32\}\times 1024\times 0.75$.. For each context length, we include 50 test samples, each comprising 1 query and 100 candidate documents.  [^4] The original version of personalized passkey retrieval uses different candidate documents for each query, resulting in 50 queries and 5,000 documents to encode for each context length. To speed up evaluation, we share the candidate documents for different queries within each context length. In this way, we can measure the effective context size of embedding models for up to 32k tokens. Examples for both synthetic tasks are presented in Figure 3. For the passkey test, the <prefix / suffix> are repeats of "The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again." For the needle test, the <prefix> and <suffix> form a long essay.
3.3 Real Tasks in LongEmbed
While synthetic tasks offer flexibility in manipulating context length and distributing key information, they still differ from real-world scenarios. To conduct a comprehensive evaluation, we have tailored following long-form QA and summarization tasks for long context retrieval. Note that for QA and summarization datasets, we use the questions and summaries as queries, respectively.
NarrativeQA (Kočiský et al., 2018) is a QA dataset comprising long stories averaging 50,474 words and corresponding questions about specific content such as characters, events. As these details are dispersed throughout the story, models must process the entire long context to get correct answers.
Table LABEL:tab:statistics presents the overall statistics of LongEmbed. Considering the computational complexity that increases quadratically with input length, we intentionally restrict the number of candidate documents in each task to to not exceed $10^{3}$. In this way, we can efficiently evaluate the basic long context capabilities of embedding models. For further elaboration on the source and examples for each dataset, please refer to Appendix C.
@@ 4 Methodology
4.1 Absolute Position Embedding (APE) & Rotary Position Embedding (RoPE)
@@
Rotary Position Embedding (RoPE) is the most pervasive position embedding strategy in the era of LLMs, including LLaMA (Touvron et al., 2023), Gemma (Team et al., 2024), QWen (Bai et al., 2023a), etc. It encodes position information of tokens with a rotation matrix that naturally incorporates explicit relative position dependency. To elucidate, given a hidden vector ${\bm{h}}=[h_{0},h_{1},...,h_{d-1}]$ of dimension $d$, and a position index $m$, RoPE operates as follows:
$f({\bm{h}},m)=[(h_{0}+\mathrm{i}h_{1})e^{\mathrm{i}m\theta_{0}},(h_{2}+\mathrm{i}h_{3})e^{\mathrm{i}m\theta_{1}},...,(h_{d-2}+\mathrm{i}h_{d-1})e^{\mathrm{i}m\theta_{d/2-1}}]$ (1)
where $\theta_{j}=10000^{-2j/d},j\in\{0,1,...,d/2-1\}$, $\mathrm{i}=\sqrt{-1}$ is the imaginary unit. Unlike APE that is directly applied to the input vector ${\bm{x}}$, RoPE is employed on the query and key vectors at each layer. The attention score $a({\bm{q}},{\bm{k}})$ between a query ${\bm{q}}$ at position $m$ and a key ${\bm{k}}$ at position $n$ is defined as:
$a({\bm{q}},{\bm{k}})$ $=\mathrm{Re}\langle f({\bm{q}},m),f({\bm{k}},n)\rangle=\mathrm{Re}\left[\sum_{j=0}^{d/2-1}(q_{2j}+\mathrm{i}q_{2j+1})(k_{2}j-\mathrm{i}k_{2j+1})e^{\mathrm{i}(m-n)\theta_{j}}\right]$
$:=g({\bm{q}},{\bm{k}},(m-n){\bm{\theta}})$ (2)
@@
4.2 Context Window Extension for APE-based Models

### v3

@@ Abstract
Embedding models play a pivotal role in modern NLP applications such as document retrieval. However, existing embedding models are limited to encoding short documents of typically 512 tokens, restrained from application scenarios requiring long inputs. This paper explores context window extension of existing embedding models, pushing their input length to a maximum of 32,768. We begin by evaluating the performance of existing embedding models using our newly constructed LongEmbed benchmark, which includes two synthetic and four real-world tasks, featuring documents of varying lengths and dispersed target information. The benchmarking results highlight huge opportunities for enhancement in current models. Via comprehensive experiments, we demonstrate that training-free context window extension strategies can effectively increase the input length of these models by several folds. Moreover, comparison of models using Absolute Position Encoding (APE) and Rotary Position Encoding (RoPE) reveals the superiority of RoPE-based embedding models in context window extension, offering empirical guidance for future models. Our benchmark, code and trained models will be released to advance the research in long context embedding models.
LongEmbed: Extending Embedding Models for Long Context Retrieval
Dawei Zhu thanks:  Contribution during Dawei’s internship at MSR Asia. Sujian Li is the corresponding author.♡♠  Liang Wang   Nan Yang   Yifan Song ♡♠  Wenhao Wu ♡♠ Furu Wei   Sujian Li ♡♠♣  School of Computer Science, Peking University  National Key Laboratory for Multimedia Information Processing, Peking University  Jiangsu Collaborative Innovation Center for Language Ability, Jiangsu Normal University  Microsoft Corporation {dwzhu,lisujian}@pku.edu.cnwangliang@microsoft.com https://github.com/dwzhu-pku/LongEmbed
@@ 1 Introduction
Text embeddings are vector representations of natural language that encode its semantic information. They play a pivotal role in various natural language processing (NLP) tasks, including information retrieval (IR) and retrieval-augmented generation (RAG). However, embedding models for producing these vector representations still operates within a very narrow context window, many supporting only 512 input tokens (Wang et al., 2022; Xiao et al., 2023; Ni et al., 2022). This narrow context window has greatly hindered their application in scenarios requiring long inputs, such as long Wikipedia articles and meeting scripts (Saad-Falcon et al., 2024).
@@
Figure 1: (a) Overview of the LongEmbed benchmark. (b) Performance of current embedding models on passkey retrieval, with evaluation length ranging from 256 to 32,768  [^2] For simplicity, we report results from the base versions of the included models by default. The supported context length of each model is presented in Table 2. Inputs exceeding the supported context length are truncated.. $\blacktriangle$ / $\blacklozenge$ denotes embedding models with 512 / $\geq$ 4k context. The greener   a cell is, the higher retrieval accuracy this model achieves on the corresponding evaluation length. (c) Effects of context window extension methods on E5, E5-RoPE, E5-Mistral, measured by improvements of Avg. Scores on LongEmbed. SE / NTK is short for SelfExtend / NTK-Aware Interpolation.
@@
For models utilizing RoPE (Su et al., 2021), substantial enhancements on LongEmbed are observed when employing methods that fully leverage RoPE’s advantages, such as NTK (Peng and Quesnelle, 2023) and SelfExtend (Jin et al., 2024). As illustrated in Figure 1(b) and 1(c), leveraging NTK extends the context window of E5-Mistral to 32k, achieving close-to-perfect accuracy on passkey retrieval and state-of-the-art performance on LongEmbed. Further, for fair comparison of APE / RoPE-based embedding models, we pre-train E5-RoPE following the training procedure and data of E5. Thorough comparison of E5 and E5-RoPE reveals the superiority of RoPE-based embedding models in context window extension. To sum up, our contributions are as follows:
  • We construct LongEmbed to benchmark long context retrieval, which includes two synthetic and four real-world tasks, featuring documents of varying lengths and dispersed target information.
  • We have conducted comprehensive experiments on training-free context window extension, extending the input length of existing embedding models by several folds.
  • We reveal the superiority of RoPE-based embedding models in context window extension via thorough comparison of models adopting APE and RoPE, offering empirical guidance for future embedding models.
  • Our benchmark and trained models (E5Base-4k, E5-RoPEBase) will be released to advance the research in long context embedding models.
@@ 2 Related Work
Text Embedding Models. Text embeddings encode semantic information of text as low-dimensional vectors, enabling numerous NLP applications. Early attempts on embeddings models include latent semantic indexing (Deerwester et al., 1990) and weighted average of word embeddings (Mikolov et al., 2013). Modern embedding models (Wang et al., 2022; Xiao et al., 2023; Neelakantan et al., 2022) exploit supervision from labeled query-document pairs, adopting a multi-stage training paradigm that pre-trained on large-scale raw text pairs using contrastive loss, then fine-tuned on small scale but high-quality datasets.
@@
Context Window Extension for LLMs. Due to the high cost of pre-training an LLM from scratch, there have been many efforts towards extending the context window of existing LLMs in a plug-and-play manner. We categorize these efforts as follows: 1) Divide-and-conquer, which involves segmenting long inputs into short chunks, processing each chunk with the model, and aggregating the results, as demonstrated by PCW (Ratner et al., 2023); 2) Position reorganization, which reorganizes position ids to accommodate longer inputs, as exemplified by SelfExtend (Jin et al., 2024), DCA (An et al., 2024). 3) Position interpolation, which introduces new position embeddings by interpolating existing ones, includes PI (Chen et al., 2023), NTK (Peng and Quesnelle, 2023), YaRN (Peng et al., 2023), and Resonance RoPE (Wang et al., 2024a). Our paper thoroughly investigates these three lines of methods on embedding models. We also acknowledge other efforts in context extension, such as token compression (Jiang et al., 2023; Ge et al., 2023; Zhang et al., 2024a) and memory-based transformers (Wang et al., 2024b; Xiao et al., 2024). However, the former is not applicable for bidirectional attention, and the latter requires complex mechanisms for accessing encoded content, hence we do not experiment with these two categories.
In addition to their plug-and-play usability, further fine-tuning on top of these methods with long training samples has been proven to yield better performance  (Xiong et al., 2023; Fu et al., 2024; Zhang et al., 2024b; Yen et al., 2024). Addressing the overhead of training on long inputs and the scarcity of extremely long training data, a line of research investigates simulating long inputs within short context, including Randomized Positions (Ruoss et al., 2023), Positional Skip-wise (PoSE) training (Zhu et al., 2023), and SkipAlign Wu et al. (2024). This paper also leverage these efforts to synthesize long training samples from the original training data, facilitating further fine-tuning on top of plug-and-play methods.
@@ 3 The LongEmbed benchmark
In this section, we first identify two limitations of existing retrieval benchmarks for evaluating long-context capabilities (§ 3.1). Then, we introduce the retrieval tasks adopted in our LongEmbed, including both synthetic ones (§ 3.2) and real ones (§ 3.3).
3.1 Examing Existing Retrieval Benchmarks
There are two main desiderata for curating a long context retrieval benchmark. First, the candidate documents should be long enough. Second, the target information to answer user query should be as uniformly distributed across the document as possible. This prevents embedding models from solely focusing on specific parts, such as the beginning (Coelho et al., 2024), to achieve unreasonably high scores. Based on these criteria, we examine existing retrieval benchmarks as follows:
Figure 2: Results of E5Base on 8 LoCo tasks that are publicly available.
BEIR Benchmark (Thakur et al., 2021) is a collection of 18 information retrieval datasets, ranging across ad-hoc web search, question answering, fact verification, etc. However, documents in this benchmark contains fewer than 300 words on average (See Table 5 in Appendix), making it unsuitable for measuring long context retrieval that usually involves documents of thousands or tens of thousands of words.
LoCo Benchmark (Saad-Falcon et al., 2024) consists 12 retrieval tasks that requires long context reasoning, spanning diverse domains such as law and finance. However, it still suffers from biased distribution of key information, as demonstrated in Figure 2. With only 512 context length, E5Base achieves >85% nDCG scores on 3 out of 8 publicly-available LoCo tasks. This severely biased distribution of target information undermines its ability to reflect model performance as context increases.
3.2 Synthetic Tasks in LongEmbed
First, we introduce the passkey and needle retrieval task for embedding models as follows:
Personalized Passkey Retrieval. Passkey retrieval (Mohtashami and Jaggi, 2023) requires LLMs to recover a random passkey hidden within a long document comprising garbage information. For embedding models, we adopt the personalized passkey retrieval (Wang et al., 2023b), where each document contains a unique person name and his/her passkey at random position. The goal is to retrieve the document containing the given person’s passkey from all candidates documents.
Figure 3: Example for the passkey and needle test. For the passkey test, the <prefix / suffix> are repeats of "The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again." For the needle test, the <prefix> and <suffix> form a long essay.
Needle-in-a-haystack Retrieval. While passkey retrieval surrounds key information with garbage sentences, needle-in-a-haystack retrieval (Kamradt, 2023; Liu et al., 2024) randomly inserts key information into an arbitrary position of a long essay, making the task more challenging. To tailor this task for embedding models, we instruct GPT-4 to generate 100 facts covering a variety of domains including physics, history, geometry, art, etc, and 100 queries correspondingly. The facts are subsequently treated as needles and randomly inserted into the PaulGrahamEssay to form 100 candidate documents. Our task is to correctly retrieve the document that contains corresponding needle given the query.
The advantage of synthetic data is that we can flexibly control context length and distribution of target information. For both tasks, we evaluate a broad context range of $\{0.25,0.5,1,2,4,8,16,32\}\times 1,024$ tokens  [^3] Since token numbers vary w.r.t. tokenizers, we use a rough estimation that 1 token = 0.75 word, and constraint the word numbers to not exceed $\{0.25,0.5,1,2,4,8,16,32\}\times 1,024\times 0.75$.. For each context length, we include 50 test samples, each comprising 1 query and 100 candidate documents.  [^4] The original version of personalized passkey retrieval uses different candidate documents for each query, resulting in 50 queries and 5,000 documents to encode for each context length. To speed up evaluation, we share the candidate documents for different queries within each context length. In this way, we can measure the effective context size of embedding models for up to 32k tokens. Examples for both tasks are in Figure 3.
3.3 Real Tasks in LongEmbed
While synthetic tasks offer flexibility in manipulating context length and distributing target information, they still differ from real-world scenarios. To conduct a comprehensive evaluation, we have tailored following long-form QA and summarization tasks for long context retrieval. For QA datasets, we use the questions as queries, the set of all input documents as candidate documents. For summarization datasets, we use the summaries as queries, and the set of all input documents as candidate documents.
NarrativeQA (Kočiský et al., 2018) is a QA dataset comprising long stories and corresponding questions about specific content such as characters, events. As these details are dispersed throughout the story, models must process the entire long context to get the correct answers.
Table 1 presents the overall statistics of LongEmbed. Considering the computational complexity that increases quadratically with input length, we intentionally restrict the number of candidate documents in each task to to not exceed $10^{3}$. In this way, we can efficiently evaluate the basic long context capabilities of embedding models. For further elaboration on the source and examples for each dataset, please refer to Appendix C.
@@ 4 Methodology
4.1 Preliminary: APE & RoPE
@@
Rotary Position Embedding (RoPE) is the most pervasive position embedding strategy in the era of LLMs, including LLaMA (Touvron et al., 2023), QWen (Bai et al., 2023a), etc. It encodes position information of tokens with a rotation matrix that naturally incorporates explicit relative position dependency. To elucidate, given a hidden vector ${\bm{h}}=[h_{0},h_{1},...,h_{d-1}]$ of dimension $d$, and a position index $m$, RoPE operates as follows:
$f({\bm{h}},m)$ $=[(h_{0}+\mathrm{i}h_{1})e^{\mathrm{i}m\theta_{0}},(h_{2}+\mathrm{i}h_{3})e^{\mathrm{i}m\theta_{1}},...,$
$(h_{d-2}+\mathrm{i}h_{d-1})e^{\mathrm{i}m\theta_{d/2-1}}]$
where $\theta_{j}=10000^{-2j/d},j\in\{0,1,...,d/2-1\}$, $\mathrm{i}=\sqrt{-1}$ is the imaginary unit. Unlike APE that is directly applied to the input vector ${\bm{x}}$, RoPE is employed on the query and key vectors at each layer. The attention score $a({\bm{q}},{\bm{k}})$ between a query ${\bm{q}}$ at position $m$ and a key ${\bm{k}}$ at position $n$ is:
[…]
@@
4.2 Extending APE-based Models